the lenght of a rectangle is 5 inches more than the width. the perimeter is 70 inches. find the lenght
Answer by dlam5(15)

represent variables for problem.
l-- length
w---width
so,
if IS means =
l=5+w
a perim. is 2l +2w
so, you now have a SYSTEM!
l=5+w
2l+2w=70
2(5+w) +2w =70
10 + 2w + 2w =70
10+ 4w =70
4w=60
w=15

If width is 15, and length is 5 more than width, l=20.
Check:
2(20) + 2(15) = 40+30 =70 :)